# Dimension conservation of harmonic measures in products of hyperbolic spaces

### v1

@@ 2.3. Random walks
The group ${\bm{\Gamma}}$ acts on $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ through $\operatorname{Isom}{\mathcal{X}}\times\operatorname{Isom}{{\mathcal{X}}^{\star}}$. It satisfies that $\nu_{\pi}$ is $\pi$-stationary, i.e.,
@@
Let us define the metric in ${\mathcal{X}}\times{{\mathcal{X}}^{\star}}$ by
@@ 3.1. Upper bounds on dimensions of conditional measures
By (3.4), the bounded martingale convergence theorem shows that for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{{\mathcal{X}}^{\star}}$,
@@
By (3.5), the bounded martingale convergence theorem shows that for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{{\mathcal{X}}^{\star}}$,
@@
$\limsup_{n\to\infty}\frac{\log\nu_{\pi}^{\eta}(\bm{B}(\bm{z}_{\infty},r))}{\log r}\leq\frac{\overline{h}+\varepsilon}{l-\varepsilon}.$ (3.10)
@@ 3.2. Lower bounds on dimensions of conditional measures
In fact, in the above the sequence $r_{n}:=e^{-ln}$ for $n\in{\mathbb{Z}}_{+}$ is replaced by positive reals $r$ tending to 00 since $r_{n+1}=e^{-l}r_{n}$ for all $n\in{\mathbb{Z}}_{+}$. Applying to Lemma 2.2 the measures $\nu_{\pi}^{\eta}$ and $F$ implies the following: There exists a constant $L\geq 1$ such that for $\nu_{{\mu^{\star}}}$-almost every $\eta\in D$, for $\nu_{\pi}^{\eta}$-almost every $(\xi,\eta)\in F$ and for a constant $r(\xi,\eta)>0$,
@@ 4.2. Lower bounds on dimensions of harmonic measures in product spaces
Let ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ where $\Gamma$ and Γsuperscriptnormal-${\Gamma^{\star}}$ are countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ with finite exponential growth relative to $({\mathcal{X}},d)$ and to $({{\mathcal{X}}^{\star}},d^{\star})$ respectively. For every probability measure $\pi$ on ${\bm{\Gamma}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$, it holds that for $\nu_{\pi}$-almost every ${\bm{\xi}}\in\partial{\mathcal{X}}\times{{\mathcal{X}}^{\star}}$,
@@
Note that ${\bm{\Gamma}}$ has a finite exponential growth relative to $({\mathcal{X}}\times{{\mathcal{X}}^{\star}},\bm{d})$ by the assumption. By Lemma 3.3, for every $\varepsilon>0$ there exists an $N\in{\mathbb{Z}}_{+}$, a Borel set $D$ in $\partial{{\mathcal{X}}^{\star}}$ with $\nu_{\mu^{\star}}(D)\geq 1-\varepsilon$ and a Borel set $F$ in $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ with $\nu_{\pi}(F)\geq 1-\varepsilon$ as stated. If we define $\hat{F}:=(\partial{\mathcal{X}}\times D)\cap F$, then
@@
Let $({\mathcal{X}},d)$ and $({{\mathcal{X}}^{\star}},d^{\star})$ be roughly geodesic hyperbolic metric spaces with bounded growth at some scale. Let ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ where $\Gamma$ and Γsuperscriptnormal-${\Gamma^{\star}}$ are countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ with finite exponential growth relative to $({\mathcal{X}},d)$ and to $({{\mathcal{X}}^{\star}},d^{\star})$ respectively. For every probability measure $\pi$ on ${\bm{\Gamma}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$, the harmonic measure $\nu_{\pi}$ on $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ is exact dimensional. Moreover, if $l({\mathcal{X}},\mu)\geq l({{\mathcal{X}}^{\star}},{\mu^{\star}})$, then it holds that
@@ 5.1. Pivotal times
A sequence of pivotal times $P_{n}\subseteq\{1,\dots,n\}$ is defined inductively as in the following: Let $P_{0}:=\emptyset$ (the empty set). Given $P_{n-1}$, let $k=0$ and $y_{k}:=o$ if $P_{n-1}=\emptyset$. Suppose that $P_{n-1}\neq\emptyset$. Let us say that the local geodesic condition is satisfied at time $n$ if
@@ 5.2. A lower bound for entropy
Let $\Gamma$ and Γsuperscriptnormal-${\Gamma^{\star}}$ be countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ respectively, and ${\bm{\Gamma}}:=\Gamma\times{\Gamma^{\star}}$. Further let us consider a probability measure $\pi$ on ${\bm{\Gamma}}$ of the following form:
@@
for some $\alpha\in(0,1]$, non-elementary probability measures $\lambda$ and ${\lambda^{\star}}$ on $\Gamma$ and on Γsuperscriptnormal-${\Gamma^{\star}}$ respectively, and a probability measure $\pi_{0}$ on ${\bm{\Gamma}}$. For the marginal ${\mu^{\star}}$ of $\pi$ on Γsuperscriptnormal-${\Gamma^{\star}}$, it holds that $h(\pi)-h(\partial{{\mathcal{X}}^{\star}},{\mu^{\star}})>0$. Moreover, if in addition ${\bm{\Gamma}}$ has a finite exponential growth relative to $({\mathcal{X}},\bm{d})$, then for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{{\mathcal{X}}^{\star}}$, the Hausdorff dimension of the conditional measure $\nu_{\pi}^{\eta}$ is positive.
@@
Conditioning on ${\mathcal{G}}$ amounts to fix a typical sequence $\{\varepsilon_{i}\}_{i=1}^{\infty}$, $\{t_{i}\}_{i=1}^{\infty}$ and a trajectory of ${\mu^{\star}}$-random walk $\{w^{\star}_{n}\}_{n\in{\mathbb{Z}}_{+}}$, increments $x_{i}$ of $\mu$-random walk outside the time intervals $(N(t_{j}-1),Nt_{j}]$ for $j=1,\dots,\tau(n)$. Under this conditioning, $s^{\prime}_{1},\dots,s^{\prime}_{\tau}$ is a sequence of independent random elements in $\Gamma$ with the common distribution $\lambda_{\mathcal{S}}^{\ast 2}$.
@@
Acknowledgment
The author is partially supported by JSPS Grant-in-Aid for Scientific Research JP20K03602.

### v2

@@ 2.3. Random walks
The group $\operatorname{Isom}{\mathcal{X}}\times\operatorname{Isom}{{\mathcal{X}}^{\star}}$ acts on $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ coordinatewise and ${\bm{\Gamma}}$ acts on $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$. It satisfies that $\nu_{\pi}$ is $\pi$-stationary, i.e.,
@@
Let us define the metric on ${\mathcal{X}}\times{{\mathcal{X}}^{\star}}$ by
@@ 3.1. Upper bounds on dimensions of conditional measures
By (3.4), the Lévy downward theorem applied to the right hand side above shows that for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{{\mathcal{X}}^{\star}}$,
@@
By (3.5), the Lévy upward theorem shows that for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{{\mathcal{X}}^{\star}}$,
@@
$\limsup_{r\to 0}\frac{\log\nu_{\pi}^{\eta}(\bm{B}(\bm{z}_{\infty},r))}{\log r}\leq\frac{\overline{h}+\varepsilon}{l-\varepsilon}.$ (3.10)
@@ 3.2. Lower bounds on dimensions of conditional measures
In fact, in the above the sequence $r_{n}:=e^{-ln}$ for $n\in{\mathbb{Z}}_{+}$ is replaced by positive reals $r$ tending to 00 since $r_{n+1}=e^{-l}r_{n}$ for all $n\in{\mathbb{Z}}_{+}$. Applying Lemma 2.2 to the measures $\nu_{\pi}^{\eta}$ and $F$ implies the following: There exists a constant $L\geq 1$ such that for $\nu_{{\mu^{\star}}}$-almost every $\eta\in D$, for $\nu_{\pi}^{\eta}$-almost every $(\xi,\eta)\in F$ and for a constant $r(\xi,\eta)>0$,
@@ 4.2. Lower bounds on dimensions of harmonic measures in product spaces
Let ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ where $\Gamma$ and ${\Gamma^{\star}}$ are countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ with finite exponential growth relative to $({\mathcal{X}},d)$ and to $({{\mathcal{X}}^{\star}},d^{\star})$ respectively. For every probability measure $\pi$ on ${\bm{\Gamma}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$, it holds that for $\nu_{\pi}$-almost every ${\bm{\xi}}\in\partial{\mathcal{X}}\times{{\mathcal{X}}^{\star}}$,
@@
Note that ${\bm{\Gamma}}$ has a finite exponential growth relative to $({\mathcal{X}}\times{{\mathcal{X}}^{\star}},\bm{d})$ by the assumption. By Lemma 3.3, for every $\varepsilon>0$ there exist an $N\in{\mathbb{Z}}_{+}$, a Borel set $D$ in $\partial{{\mathcal{X}}^{\star}}$ with $\nu_{\mu^{\star}}(D)\geq 1-\varepsilon$ and a Borel set $F$ in $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ with $\nu_{\pi}(F)\geq 1-\varepsilon$ as stated. If we define $\hat{F}:=(\partial{\mathcal{X}}\times D)\cap F$, then
@@
Let $({\mathcal{X}},d)$ and $({{\mathcal{X}}^{\star}},d^{\star})$ be roughly geodesic hyperbolic metric spaces with bounded growth at some scale. Let ${\bm{\Gamma}}=\Gamma\times{\Gamma^{\star}}$ where $\Gamma$ and ${\Gamma^{\star}}$ are countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ with finite exponential growth relative to $({\mathcal{X}},d)$ and to $({{\mathcal{X}}^{\star}},d^{\star})$ respectively. For every probability measure $\pi$ on ${\bm{\Gamma}}$ with finite first moment and non-elementary marginals $\mu$ and ${\mu^{\star}}$, the harmonic measure $\nu_{\pi}$ on $\partial{\mathcal{X}}\times\partial{{\mathcal{X}}^{\star}}$ is exact dimensional. Moreover, if $l({\mathcal{X}},\mu)\geq l({{\mathcal{X}}^{\star}},{\mu^{\star}})$, then it holds that
@@ 5.1. Pivotal times
A sequence of pivotal times $P_{n}\subseteq\{1,\dots,n\}$ is defined inductively as in the following: Let $P_{0}:=\emptyset$ (the empty set). Given $P_{n-1}$, let $k:=\max P_{n-1}$ if $P_{n-1}\neq\emptyset$, and let $k=0$ and $y_{k}:=o$ if $P_{n-1}=\emptyset$. Suppose that $P_{n-1}\neq\emptyset$. Let us say that the local geodesic condition is satisfied at time $n$ if
@@ 5.2. A lower bound for entropy
Let $\Gamma$ and ${\Gamma^{\star}}$ be countable subgroups in $\operatorname{Isom}{\mathcal{X}}$ and in $\operatorname{Isom}{{\mathcal{X}}^{\star}}$ respectively, and ${\bm{\Gamma}}:=\Gamma\times{\Gamma^{\star}}$. Further let us consider a probability measure $\pi$ on ${\bm{\Gamma}}$ of the following form:
@@
for some $\alpha\in(0,1]$, non-elementary probability measures $\lambda$ and ${\lambda^{\star}}$ on $\Gamma$ and on ${\Gamma^{\star}}$ respectively, and a probability measure $\pi_{0}$ on ${\bm{\Gamma}}$. Then for the marginal ${\mu^{\star}}$ of $\pi$ on ${\Gamma^{\star}}$, it holds that $h(\pi)-h(\partial{{\mathcal{X}}^{\star}},{\mu^{\star}})>0$. Moreover, if in addition ${\bm{\Gamma}}$ has a finite exponential growth relative to $({\mathcal{X}},\bm{d})$, then for $\nu_{{\mu^{\star}}}$-almost every $\eta\in\partial{{\mathcal{X}}^{\star}}$, the Hausdorff dimension of the conditional measure $\nu_{\pi}^{\eta}$ is positive.
@@
Conditioning on ${\mathcal{G}}$ amounts to fixing a typical sequence $\{\varepsilon_{i}\}_{i=1}^{\infty}$, $\{t_{i}\}_{i=1}^{\infty}$ and a trajectory of ${\mu^{\star}}$-random walk $\{w^{\star}_{n}\}_{n\in{\mathbb{Z}}_{+}}$, increments $x_{i}$ of $\mu$-random walk outside the time intervals $(N(t_{j}-1),Nt_{j}]$ for $j=1,\dots,\tau(n)$. Under this conditioning, $s^{\prime}_{1},\dots,s^{\prime}_{\tau}$ is a sequence of independent random elements in $\Gamma$ with the common distribution $\lambda_{\mathcal{S}}^{\ast 2}$.
@@
Acknowledgments
The author would like to thank the anonymous referee for a meticulous review and suggestions. The author is partially supported by JSPS Grant-in-Aid for Scientific Research JP20K03602 and JP24K06711.